OpenStudy (anonymous):

15 people volunteer for a park cleanup. the number of volunteers increases by 7 people each month for several months. After how many months will there be 50 volunteers?

OpenStudy (amistre64):

an=15+7(n-1)

OpenStudy (amistre64):

50 = 15+7n-7 50 = 8+7n 42 = 7n 6=n